# Uncovering Reliable Indicators: Improving IoC Extraction from Threat Reports

### v1

@@ Abstract
Indicators of Compromise (IoCs) are critical for threat detection and response, marking malicious activity across networks and systems. Yet, the effectiveness of automated IoC extraction systems is fundamentally limited by one key issue: the lack of high-quality ground truth. Current extraction tools rely either on manually extracted ground truth, which is labor-intensive and costly, or on automated ground truth creation methods that include non-malicious artifacts, leading to inflated false positive (FP) rates and unreliable threat intelligence. In this work, we analyze the shortcomings of existing ground truth creation strategies and address them by introducing the first hybrid human-in-the-loop pipeline for IoC extraction, which combines a large language model–based classifier (LANCE) with expert analyst validation. Our system improves precision through explainable, context-aware labeling and reduces analysts’ work factor by 43% compared to manual annotation, as demonstrated in our evaluation with six analysts. Using this approach, we produce PRISM, a high-quality, publicly available benchmark of 1,791 labeled IoCs from 50 real-world threat reports. PRISM supports both fair evaluation and training of IoC extraction methods and enables reproducible research grounded in expert-validated indicators.
@@ 1 Introduction
Detecting and analyzing malicious activity is essential in modern cybersecurity defense. Indicators of Compromise (IoCs), such as IP addresses, domain names, URLs, and file hashes, which signal malicious activity, are among the most important parts of threat intelligence [1, 2, 3]. Analysts rely on IoCs for both immediate defense, such as creating block lists and firewall rules, and for activities such as threat hunting. Accurate and timely identification of IoCs can therefore significantly improve an organization’s ability to detect, respond to, and mitigate cyber-threats.
@@
One of the most prominent sources of IoCs is threat reports (TRs), which are published by cybersecurity companies, research institutions, and government agencies, and contain critical information about attacks, campaigns, and incidents [4, 2, 5, 1]. Threat reports vary widely in structure and detail due to the lack of standardized formatting. This inconsistency makes accurate IoC extraction challenging, as indicators may be embedded in diverse formats and contexts, yet their precise identification is critical for effective threat response.
@@
Previous work has focused on three general approaches for extracting IoCs from threat reports: manual labeling, automated extraction, and automated retrieval from threat intelligence sharing platforms. As noted repeatedly [6, 1, 7, 8], manual labeling is the most accurate approach to extract IoCs from unstructured threat reports. However, this requires significant domain knowledge and can be very time-consuming, making it expensive and difficult to scale [1, 6, 7]. Automated IoC extraction tools, including recent work that leverage large language models (LLMs) [9, 10, 11], offer potential efficiency but often lack the necessary accuracy due to the complexity of natural language in technical cybersecurity reports, the variability in IoC presentation, and the difficulty in distinguishing true IoCs from similar-looking but benign artifacts (e.g., indicators such as benign domain names, IP addresses, and URLs that are not related to the threats described in the reports). Finally, while there exist efforts to distribute IoCs to the community in an easy-to-parse structured format, for instance via threat exchanges (e.g., AlienVault [12]), threat report databases (e.g., orkl.eu), and online security scanners (e.g., VirusTotal [13, 14, 15, 16]), we find that popular threat report IoC exchanges like AlienVault, often automatically extract IoCs from unstructured TRs with low precision (<<<80%), recall (<<<80%) and/or coverage. As a result, these approaches either incur high manual work factor costs or yield unreliable intelligence, underscoring the need for a new solution that is both accurate and efficient.
@@
To address these challenges, we propose the first IoC extraction approach that combines automated extraction supported by explainable AI with human-in-the-loop verification. Our central hypothesis is that large language models (LLMs), when guided by properly engineered prompts and paired with analyst feedback, can match or surpass the accuracy of manual-only labeling while significantly reducing the time cost of human effort. In our evaluation, we observe a 43% reduction in parsing time compared to manual-only workflows. Importantly, our pipeline design and prompt engineering strategies are model-agnostic and generalize across most state-of-the-art LLMs, enabling flexible deployment in diverse environments. By integrating human analysts in the IoC extraction process, we can leverage the speed of automation while ensuring high precision through human judgment. Similar “human-in-the-loop” (HITL) approaches have been successfully demonstrated in other domains where high-accuracy data labeling and expert oversight are critical, such as medical applications and some computer vision and natural language processing tasks [17, 18, 19, 20]. To the best of our knowledge, we are the first to adapt the HITL approach for the efficient and accurate labeling of IoCs from unstructured reports. To this end, we develop a HITL system that provides: (1) visual assistance (via highlighted text) that allows analysts to quickly identify indicators of interest in unstructured (potentially long) threat reports; and (2) color-coded IoC labels assigned by an LLM, each accompanied by an explanation of the contextual cues used to justify the label – designed following principles of effective data visualization [21, 22, 23].
@@
  • PRISM: The first open, fully manually validated, and well-documented benchmark dataset of IoCs from threat reports, consisting of 1,791 indicators extracted from 50 real-world threat reports. Our dataset, called PRISM, is designed to support both future training/testing of IoC extraction/labeling models and as a common evaluation dataset for comparing IoC extraction systems among each other.
We will make the HITL system, LANCE, and PRISM available to the research community upon publication.
Figure 1: Overview of the human in the loop (HITL) IoC extraction pipeline. It combines automated extraction and LLM-based labeling (LANCE) with manual annotation, thus promoting efficient, high-confidence IoC labeling.
@@ 2.1 IoC Extraction Challenges and Previous Work
Extracting Indicators of Compromise (IoCs) from unstructured threat reports is far from straightforward. While IoCs like IP addresses, domains, URLs, and file hashes are vital for threat detection and response, they are often buried in reports that lack consistent formatting. These indicators may appear inline, in tables, or within images, and their relevant context may be scattered throughout the document [8]. This fragmented presentation makes it difficult for automated tools to reliably detect and interpret IoCs.
A further complication lies in differentiating true IoCs from benign artifacts that share similar structure [24, 5]. Regular expression–based extractors, though widely used, frequently overreport by flagging non-malicious entities such as popular domains embedded within malicious URLs [1]. Because maliciousness is context-dependent, identifying IoCs accurately requires understanding context that is not always located near the indicator itself, making it difficult for NLP methods to draw accurate associations.
To address these challenges, a range of extraction methods have been proposed, spanning rule-based approaches [1, 25], NLP [8, 26, 27], machine learning (ML) [28, 29, 30, 31], deep learning (DL) [7, 32, 24, 2, 6, 5, 33], including LLMs [9, 10, 11], and graph-based techniques [34, 35]. Yet each approach has its drawbacks. Rule-based systems lack contextual awareness. NLP models often struggle with inconsistent grammar, distant dependencies, and report structure [29]. ML and DL techniques require large, accurately labeled datasets, which are scarce [8, 6, 2]. LLM-based approaches are prone to hallucinations and are limited by their context window [9, 10, 11, 36, 37].
@@ 2.2 IoC Ground Truth Generation Methods
Incorrect ground truth (GT) data, with mislabeled benign/malicious indicators, harms both training and evaluation of automated systems [38, 39]. Models trained on noisy labels risk learning incorrect patterns, overfitting, or failing to generalize. For example, labeling “github.com” as malicious teaches the model to overreport benign indicators, while missing actual threats reduces future detection capability. Ambiguity in what constitutes maliciousness, often context-dependent, further complicates matters.
@@
VirusTotal-Based Validation: More than one third of the works, mainly ML systems [28, 29, 30, 31] but also DL models [24] and graph-based implementations [35], use VirusTotal (VT) to validate extracted IoCs. While VT offers broad coverage, threshold choice is inconsistent across studies, and VT can introduce false positives or negatives [14, 15, 16].
@@
[…]
@@ 2.3 Investigating Errors of Existing Ground Truth Creation Systems
To understand the errors that common ground truth creation methods make and their sources, we review two representative systems that are often used to collect IoCs found in unstructured threat reports, IoC Searcher [1] and AlienVault [7, 12]. We evaluate the tools on 10 threat reports published in 2024, sourced from AlienVault’s database, and parsed manually by two analysts, resulting in 517 confirmed IoCs. IoC Searcher is a state-of-the-art rule-based tool with public code that has been compared to other similar tools by previous work [1, 30], and this evaluation provides a robust baseline for assessing both tools.
IoC Searcher produced 126 false positives. The most common error was “Not Malicious” (52 instances), where benign entities like domains (e.g., “msn.com”, “github.com”), IPs, (e.g., “1.1.1.1”, “4.2.2.4”), URLs, and hashes were misclassified as IoCs solely based on their format, without considering the surrounding context indicating they were benign. A second major error type involving semantic misclassification (“File Path”), was benign paths flagged as domains/URLs due to superficial string patterns (e.g., “do.zip”, “asp.net”). Other types of errors occurred by “TLD” misclassified as IoC domains, indicators pointing to the “Reporting Company” extracted as IoCs, substrings of an actual IoC, and indicators with placeholders or reduced parts in them (e.g., “www.[reducted].com”, “microsoft-update-com.github.io/XXX/update.html?id=[GUID”). These results reflect the limits in contextual understanding and formatting variability. Despite these issues, IoC Searcher showed high recall, missing only 9 indicators. Most errors stemmed from newlines splitting indicators, incorrect type detection, or partial matches (“Found Longer Variant”).
@@ 3.1 Problem Definition, Goals, and Constraints
In Section 2.2 we examined the ground truth creation methodology of 18 prominent IoC extraction papers, analyzing the transparency of their methods, the availability, and usability of their code and datasets, as summarized in Table I. We observed significant variance in how ground truth was created. Manual labeling was often poorly documented, making it unclear what role human input played in the dataset creation process. Furthermore, only one paper made both its code and dataset publicly available, with one more sharing them upon request. These limitations have been echoed in recent studies [43, 1, 2, 8, 4].
@@
To overcome these limitations, we introduce a novel IoC extraction and labeling methodology based on a Human-in-the-Loop (HITL) framework, which we use to construct the first openly available and thoroughly documented benchmark dataset for IoC extraction from threat reports, called PRISM. Our HITL pipeline integrates large language model (LLM)–generated labels with analyst oversight, enabling both high-precision labeling and significant reductions in human effort. Rather than treating human labeling and automation as mutually exclusive approaches, we combine them into a cohesive workflow in which explainable LLM predictions guide human decisions, and human validation resolves ambiguity.
@@ 3.2 LANCE
To streamline the IoC extraction and labeling process, so we can minimize the time the analysts need to spend on the task, we developed LANCE: LLM-Assisted Notation and Classification Engine, a tool that automatically extracts and labels Indicators of Compromise (IoCs) from unstructured threat reports. An overview of LANCE can be seen in Figure 2. The extraction phase utilizes a rule-based approach using regular expressions, while the labeling is performed by a Large Language Model (LLM). Explainability is central to our HITL design. By requiring the LLM to justify each classification decision, we shift the downstream analyst’s role from being a manual extractor to that of an informed validator.
@@ 3.2.1 Rule-based Extraction
Regular expression-based tools are highly effective in achieving high recall for IoC extraction but struggle with precision as discussed in Section 2.3, as they cannot differentiate between malicious and non-malicious indicators. Given the above, we employed regular expressions for the extraction step to ensure broad coverage of potential IoCs.
While RegEx-based tools like IoC Searcher maximize recall, they lack contextual understanding. In our pipeline, this extraction phase is intentionally broad: it ensures full coverage so that no potential IoC is missed, with human and LLM refinement steps later correcting for over-inclusion. This supports a recall-first strategy where precision is restored through layered validation.
@@ 3.2.2 LLM-based Indicator Labeling
To handle long reports, we implemented a rolling window approach with a segment size of approximately 8000 characters (Figure 2 part 3), ensuring splits occurred at the nearest whitespace to preserve readability. This approach has been shown to improve the performance of LLM [44, 45]. To mitigate content loss at segment boundaries, we introduced a 50% overlap between consecutive segments. This implementation ensured that no critical context was lost in segmentation [46].
@@
These design elements significantly reduced error rates, particularly in complex cases, including domains and URLs, where indicators often resemble benign artifacts or contain embedded nonIoC substrings. A detailed analysis of the prompt engineering phase and evaluation of the prompts can be found in Appendix A.

### v2

# Revealing the True Indicators: Understanding and Improving IoC Extraction From Threat Reports

@@ Abstract
Indicators of Compromise (IoCs) are critical for threat detection and response, marking malicious activity across networks and systems. Yet, the effectiveness of automated IoC extraction systems is fundamentally limited by one key issue: the lack of high-quality ground truth. Current extraction tools rely either on manually extracted ground truth, which is labor-intensive and costly, or on automated ground truth creation methods that include non-malicious artifacts, leading to inflated false positive (FP) rates and unreliable threat intelligence. In this work, we analyze the shortcomings of existing ground truth creation strategies and address them by introducing the first hybrid human-in-the-loop pipeline for IoC extraction, which combines a large language model-based classifier (LANCE) with expert analyst validation. Our system improves precision through explainable, context-aware labeling and reduces analysts’ work factor by 43% compared to manual annotation, as demonstrated in our evaluation with six analysts. Using this approach, we produce PRISM, a high-quality, publicly available benchmark of 1,791 labeled IoCs from 50 real-world threat reports. PRISM supports both fair evaluation and training of IoC extraction methods and enables reproducible research grounded in expert-validated indicators.
@@ 1 Introduction
Detecting and analyzing malicious activity is essential in modern cybersecurity defense. Indicators of Compromise (IoCs), such as IP addresses, domain names, URLs, and file hashes, which signal malicious activity, are among the most important parts of threat intelligence [1, 2, 3]. Analysts rely on IoCs for both immediate defense, such as creating block lists and firewall rules, and for activities such as threat hunting. Accurate and timely identification of IoCs can therefore significantly improve an organization’s ability to detect, respond to, and mitigate cyber threats.
@@
One of the most prominent sources of IoCs is threat reports (TRs), which cybersecurity companies, research institutions, and government agencies publish to share critical information about attacks, campaigns, and incidents [4, 2, 5, 1]. Threat reports vary widely in structure and detail due to the lack of standardized formatting. This inconsistency makes accurate IoC extraction challenging, as reports may embed indicators in diverse formats and contexts, yet their precise identification is critical for effective threat response.
@@
Previous work has focused on three general approaches for extracting IoCs from threat reports: manual labeling, automated extraction, and automated retrieval from threat intelligence sharing platforms. As noted repeatedly [6, 1, 7, 8], manual labeling is the most accurate approach to extract IoCs from unstructured threat reports. However, this requires significant domain knowledge and can be very time-consuming, making it expensive and difficult to scale [1, 6, 7]. Automated IoC extraction tools, including recent work leveraging large language models (LLMs) [9, 10, 11], offer potential efficiency but often lack the necessary accuracy due to the complexity of natural language in technical cybersecurity reports, the variability in IoC presentation, and the difficulty in distinguishing true IoCs from similar-looking but benign artifacts (e.g., indicators such as benign domain names, IP addresses, and URLs that are not related to the threats described in the reports). Finally, while there exist efforts to distribute IoCs to the community in an easy-to-parse structured format, for instance via threat exchanges (e.g., AlienVault [12]), threat report databases (e.g., orkl.eu), and online security scanners (e.g., VirusTotal [13, 14, 15, 16]), we find that popular threat report IoC exchanges, like AlienVault, often automatically extract IoCs from unstructured TRs with low precision ($80\%$), recall ($80\%$), and/or coverage. As a result, these approaches either incur high manual work factor costs or yield unreliable intelligence, underscoring the need for a new solution that is both accurate and efficient.
@@
To address these challenges, we propose the first IoC extraction approach that combines automated extraction supported by explainable AI with human-in-the-loop verification. Our central hypothesis is that large language models (LLMs), when guided by properly engineered prompts and paired with analyst feedback, can match or surpass the accuracy of manual-only labeling while significantly reducing the time cost of human effort. In our evaluation, we observe a 43% reduction in parsing time compared to manual-only workflows. Importantly, our pipeline design and prompt engineering strategies are model-agnostic and generalize across most state-of-the-art LLMs, enabling flexible deployment in diverse environments. By integrating human analysts in the IoC extraction process, we can leverage the speed of automation while ensuring high precision through human judgment. Similar “human-in-the-loop” (HITL) approaches have been successfully demonstrated in other domains where high-accuracy data labeling and expert oversight are critical, such as medical applications and some computer vision and natural language processing tasks [17, 18, 19, 20]. To the best of our knowledge, we are the first to adapt the HITL approach for the efficient and accurate labeling of IoCs from unstructured reports. To this end, we develop a HITL system that provides: (1) visual assistance (via highlighted text) that allows analysts to quickly identify indicators of interest in unstructured (potentially long) threat reports; and (2) color-coded IoC labels assigned by an LLM, each accompanied by an explanation of the contextual cues used to justify the label, designed following principles of effective data visualization [21, 22, 23].
@@
  • PRISM: The first open, fully manually validated, and well-documented benchmark dataset of IoCs from threat reports, consisting of 1,791 indicators extracted from 50 real-world threat reports. Our dataset, called PRISM, is designed to both future training and testing of IoC extraction/labeling models and serve as a common evaluation dataset for comparing IoC extraction systems among each other.
The HITL system, LANCE, and PRISM are freely available at https://github.com/EvanFr/LANCE
Figure 1: Overview of the human-in-the-loop (HITL) IoC extraction pipeline. It combines automated extraction and LLM-based labeling (LANCE) with manual annotation, thus promoting efficient, high-confidence IoC labeling.
@@ 2.1 IoC Extraction Challenges and Previous Work
Extracting Indicators of Compromise (IoCs) from unstructured threat reports is far from straightforward. While IoCs like IP addresses, domains, URLs, and file hashes are vital for threat detection and response, they are often buried in reports that lack consistent formatting. These indicators may appear inline, in tables, or within images, and their relevant context may be scattered throughout the document [8]. This fragmented presentation makes it difficult for automated tools to detect and interpret IoCs reliably.
A further complication lies in differentiating true IoCs from benign artifacts that share a similar structure [24, 5]. Regular expression–based extractors, though widely used, frequently overreport by flagging non-malicious entities such as popular domains embedded within malicious URLs [1]. Because maliciousness is context-dependent, accurately identifying IoCs requires understanding context that is not always located near the indicator itself, making it difficult for NLP methods to draw accurate associations.
To address these challenges, researchers have proposed a range of extraction methods, spanning rule-based approaches [1, 25], NLP [8, 26, 27], machine learning (ML) [28, 29, 30, 31], deep learning (DL) [7, 32, 24, 2, 6, 5, 33], including LLMs [9, 10, 11], and graph-based techniques [34, 35]. However, each approach has its drawbacks. Rule-based systems lack contextual awareness. NLP models often struggle with inconsistent grammar, distant dependencies, and report structure [29]. ML and DL techniques require large, accurately labeled datasets, which are scarce [8, 6, 2]. LLM-based approaches are prone to hallucinations and are limited by their context window [9, 10, 11, 36, 37].
@@
In addition to IoC-focused extraction systems, recent works have broadened CTI extraction to include relations and attack behavior. For instance, Extractor [38], Looking Beyond IoCs (LADDER) [39], and CTINexus [40] leverage NLP and LLMs to extract complex threat intelligence beyond indicators (e.g., attack patterns and knowledge graphs). While these systems offer a broader scope, they rely heavily on RegEx for indicator extraction or do not target high-precision IoC labeling. For example, CTINexus uses subject–verb–object triplet extraction with LLMs but exhibits low recall in our replication (see Appendix B). In contrast, our work focuses on precise, context-aware IoC labeling and complements these systems by improving the fidelity of extracted indicators.
@@ 2.2 IoC Ground Truth Generation Methods
Incorrect ground truth (GT) data, with mislabeled benign/malicious indicators, harms both training and evaluation of automated systems [41, 42]. Models trained on noisy labels risk learning incorrect patterns, overfitting, or failing to generalize. For example, labeling github.com as malicious teaches the model to overreport benign indicators, while missing actual threats reduces future detection capability. Ambiguity in what constitutes maliciousness, often context-dependent, further complicates matters.
@@
VirusTotal-Based Validation: More than one-third of the works, mainly ML systems [28, 29, 30, 31] but also DL models [24] and graph-based implementations [35], use VirusTotal (VT) to validate extracted IoCs. While VT offers broad coverage, threshold choice is inconsistent across studies, and VT can introduce false positives or negatives [14, 15, 16].
@@
TABLE I: Summary of IoC extraction tools and their ground truth (GT) creation methods, grouped by extraction technique. The Manual column denotes the extent of manual labeling, with & indicating extensive and clearly documented manual labeling, partial manual labeling with documentation, + minimal manual labeling with documentation, and “–” a claim of manual labeling but without sufficient explanation. Check marks (✓) in the GT Creation Methods columns indicate that the paper employed the corresponding method. We report malicious and non-malicious samples’ GT sizes where available. ✓ in Availability indicates publicly available code and/or dataset. A & in the Usability column marks that we successfully used the tool, while a indicates that the tool was usable but out of scope (e.g., works on input types that do not include threat reports).
@@ 2.3 Investigating Errors of Existing Ground Truth Creation Systems
To understand the errors that common ground truth creation methods make and their sources, we review two representative systems that are often used to collect IoCs found in unstructured threat reports, IoC Searcher [1] and AlienVault [7, 12]. We evaluate the tools on 10 threat reports published in 2024, sourced from AlienVault’s database, and parsed manually by two analysts, resulting in 517 confirmed IoCs. IoC Searcher is a state-of-the-art rule-based tool with public code that has been compared to other similar tools by previous work [1, 30]. The following evaluation provides a robust baseline for assessing both tools.
IoC Searcher produced 126 false positives. The most common error was “Not Malicious” (52 instances), where benign entities like domains (e.g., msn.com, github.com), IPs, (e.g., 1.1.1.1, 4.2.2.4), URLs, and hashes were misclassified as IoCs solely based on their format, without considering the surrounding context indicating them as benign. Next came ”File Path,” a semantic misclassification where benign paths were flagged as domains or URLs due to shared string patterns (e.g., do.zip, asp.net). Other common errors included TLD misclassification as IoC domains, indicators pointing to the “Reporting Company” extracted as IoCs, partial substrings extraction of the full IoC, and extracting indicators with placeholders or redacted labels(e.g., www.[redacted].com, microsoft-update-com.github.io/XXX/update.html?id=[GUID ). These results reflect the limits of these tools in contextual understanding and formatting variability. Despite these issues, IoC Searcher showed high recall, missing only nine indicators. After manually investigating the misclassified IoCs, we found that most errors stemmed from newlines splitting indicators, incorrect type detection, or partial matches (“Found Longer Variant”).
@@ 3 IoC Labeling Methodology
We limited our scope to the four most common and reliably verifiable indicator types, IP addresses, domain names, URLs, and file hashes, which are well-supported by VirusTotal and commonly shared in structured platforms such as AlienVault. While other indicators (e.g., Bitcoin addresses, email addresses, file paths, registry keys) are indeed relevant, they are harder to verify at scale using current automated methods, and less consistently present across reports. We leave the inclusion of such indicators to future work.
@@ 3.1 Problem Definition, Goals, and Constraints
In Section 2.2, we examined the ground truth creation methodology of 18 prominent IoC extraction papers, analyzing the transparency of their methods, as well as the availability, and usability of their code and datasets, as summarized in Table I. We observed significant variance in how these works created ground truth. Manual labeling was often poorly documented, making it unclear what role human input played in the dataset creation process. Furthermore, only one paper made both its code and dataset publicly available, with one more sharing them upon request. These limitations have been echoed in recent studies [46, 1, 2, 8, 4].
@@
To overcome these limitations, we introduce a novel IoC extraction and labeling methodology based on a human-in-the-loop (HITL) framework. We then use to construct the first openly available and thoroughly documented benchmark dataset for IoC extraction from threat reports, called PRISM. Our HITL pipeline integrates large language model (LLM)–generated labels with analyst oversight, enabling both high-precision labeling and significant reductions in human effort. Rather than treating human labeling and automation as mutually exclusive approaches, we combine them into a cohesive workflow in which explainable LLM predictions guide human decisions, and human validation resolves ambiguity.
@@
Throughout this work, we distinguish between labeling, which refers to assigning IoC/nonIoC status to individual indicators, and ground truth generation, which refers to the full process of creating high-quality labels through labeling, human validation, and dispute resolution.
@@ 3.2 LANCE
To streamline the IoC extraction and labeling process, so we can minimize the time the analysts need to spend on the task, we developed LANCE: LLM-Assisted Notation and Classification Engine, a tool that automatically extracts and labels Indicators of Compromise (IoCs) from unstructured threat reports. An overview of LANCE can be seen in Figure 2. The extraction phase utilizes a rule-based approach using regular expressions, while a Large Language Model (LLM) performs the labeling. Explainability is central to our HITL design. By requiring the LLM to justify each classification decision, we shift the downstream analyst’s role from being a manual extractor to that of an informed validator.
@@ 3.2.1 Rule-based Extraction
Regular expression-based tools are highly effective in achieving high recall for IoC extraction, but struggle with precision. As discussed in Section 2.3, they fail to differentiate between malicious and non-malicious indicators. Given the above, we employed regular expressions for the extraction step to ensure broad coverage of potential IoCs.
While RegEx-based tools like IoC Searcher maximize recall, they lack contextual understanding. In our pipeline, this extraction phase is intentionally broad: it ensures full coverage so that no potential IoC is missed, with human and LLM refinement steps later correcting for over-inclusion. This approach supports a recall-first strategy where precision is restored through layered validation.
@@ 3.2.2 LLM-based Indicator Labeling
To handle long reports, we implemented a rolling window approach with a segment size of approximately 8000 characters (Figure 2 part 3), ensuring splits occurred at the nearest whitespace to preserve readability. This approach has been shown to improve the performance of LLM [47, 48], allowing LANCE to handle reports of arbitrary length. To mitigate content loss at segment boundaries, we introduced a 50% overlap between consecutive segments. This implementation ensured that no critical context was lost in segmentation [49].
@@
These design elements significantly reduced error rates, particularly in complex cases, including domains and URLs, where indicators often resemble benign artifacts or contain embedded nonIoC substrings. We provide a detailed analysis of the prompt engineering phase and an evaluation of the prompts in Appendix A.